Let a,b ∈ R with a < b. Let Ta,b ⊆ Q be the subset defined by Ta,b :=r ∈ Q : a<r<b. Prove that the set Ta,b is infinite.


1
Expert's answer
2021-01-19T04:01:28-0500

Let's first prove that this set is non-empty. As "a<b", there is such "n\in \mathbb{N}" that "b-a>\frac{2}{n}". Indeed, if that would not be the case, that would mean that "a-b \geq -\frac{2}{n}, \forall n\in\mathbb{N}", then by the existence of a supremum, "a-b\geq \sup\{-2/n, n\in\mathbb{N} \}=0", "a\geq b" which is a contradiction. Now as "b-a>\frac{2}{n}" for some fixed "n", there is such "m\in \mathbb{Z}" that "\frac{m-1}{n}< a\leq\frac{m}{n}" ("m=\min\{x\in\mathbb{Z}, x\geq na \}" and this set is non-empty and has a lower bound by the Archimedian property of real numbers). So "\frac{2m+1}{2n} \in \mathbb{Q}", "\frac{2m+1}{2n}>\frac{m}{n}\geq a". In addition, as "b-a>\frac{2}{n}" and "\frac{m}{n}-a<\frac{1}{n}"(by construction of "m") , we have "\frac{2m+1}{2n}<b". Therefore there is a rational number "r, a<r<b", so the set "T_{a,b}" is non-empty. But it is enough to prove that it is infinite: as there is a rational number "r" in "T_{a,b}" for any real numbers "a,b" there is a rational number "r'" in "T_{a,r}" which is also in "T_{a,b}" ("a<r'<r<b"). THus there is at least two rational numbers "a<r'<r<b". But therefore all the possible combinations of a form "q\in\mathbb{Q}, 0\leq q\leq1" , "a<r'<q\cdot r' + (1-q)\cdot r<r<b" are in "T_{a,b}". As there is infinitely many such combinations, the set "T_{a,b}" is infinite.